### v1

# An Adverse Weather-Immune Scheme with Unfolded Regularization and Foundation Model Knowledge Distillation for Street Scene Understanding

Abstract
Various adverse weather conditions pose a significant challenge to autonomous driving (AD) perception. A common strategy is to minimize the disparity between images captured in clear and adverse weather conditions. However, this technique typically relies on utilizing clear image as a reference, which is challenging to obtain in practice. Furthermore, this method typically targets a single adverse condition and perform poorly when confronting the mixup of multiple adverse weather conditions. To address these issues, we introduce a reference-free and Adverse weather-Immune scheme (called AdvImmu) achieved by leveraging the invariance of weather conditions over short periods (seconds). Specifically, AdvImmu includes three components: Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs). LSM leverages temporal correlations between adjacent frames to enhance model performance. GSM is proposed to shuffle LSM segments to prevent the overfitting to temporal patterns of only using LSM. URs are the deep unfolding implementation of two proposed regularizers to penalize the model complexity to enhance across-weather generalization. In addition, to overcome the over-reliance on consecutive frame-wise annotations in the training of AdvImmu (typically unavailable in AD scenarios), we incorporate the Segment Anything Model (SAM) to annotate frames, and additionally propose a cluster algorithm (denoted as SBICAC) to surmount SAM’s category-agnostic issue to generate pseudo-labels. Extensive experiments demonstrate that the proposed AdvImmu outperforms existing state-of-the-art methods by 88.56% in mean Intersection over Union (mIoU).
I Introduction
Autonomous driving (AD) perception is critical [1, 2, 3] and is still a significant challenge in various adverse weather conditions so far [4]. Existing domain adaption approach to address this challenge involves reducing the disparity between images captured in clear weather as a reference and those taken in adverse conditions [5, 6]. However, obtaining such a reference image can be challenging in reality. Furthermore, these methods are generally developed to handle specific adverse condition, such as foggy weather. This specialization often leads to a significant decline in performance when vehicles encounter unusual weather conditions where the model has not been specifically tuned.
To effectively tackle these challenges, we propose a novel scheme (named AdvImmu). AdvImmu benefits from being independent of reference images and is robust against various adverse weather conditions, which is materialized by considering the stability of weather conditions within past seconds. Concretely, AdvImmu consists of Locally Sequential Mechanism (LSM), Globally Shuffled Mechanism (GSM), and Unfolded Regularizers (URs) modules. Specifically, LSM exploits the temporal correlations between consecutive frames via three distinct units: Instant Unit (InsU), which is designed to detect instantaneous information in the input; Integral Unit (IntU), which focuses on capturing stable and shared background information from the scenes; Derivative Unit (DU), which excels in sensing dynamic changes in the environment. As discussed, since LSM integrates temporal patterns from historical consecutive frames, it is susceptible to overfitting specific input sequence order. To mitigate LSM’s temporality-related overfitting, GSM is introduced to rearrange the segments processed by the LSM. This strategic shuffling ensures that the model does not merely memorize a particular data order but instead learns a generalizable pattern across varied data sequences. On top of LSM and GSM, URs are proposed to improve the model’s generalization across diverse adverse weather conditions by penalizing the model complexity. Concretely, in addition to the conventional pixel-level cross entropy loss used in semantic understanding, we propose using the deep unfolding technique [7] to unfold an image-level regularizer and an inter-class contrastive regularizer into layers. Deep unfolding can avoid heuristics or exhaustive searches on the weights of both regularizers, which prevents instability and suboptimal performance. The proposed AdvImmu is summarized in Fig. 1.
Figure 1: Overview of the proposed AdvImmu. GSM shuffles and groups the LSM segments to avoid the overfitting to temporal patterns, whereas GSM-Free does not. In the context of LSM, $T$ denotes the current frame number in the input sequence, while $t$ (i.e., LSM depth) refers to the count of consecutive frames prior to frame $T$.
The training of the proposed AdvImmu requires consecutive frame-wise annotation for the images in training datasets, which is impractical in the realm of AD scenarios. To tackle this challenge, we propose to use the Segment Anything Model (SAM) [8] to annotate the images. However, since SAM-generated masks lack of category information (e.g., different trees are assigned with different semantic IDs), we additionally propose a Similarity-based Iterative Cluster Assignments and Centroids Refinement (SBICAC) algorithm to overcome SAM’s category-agnostic issue and further to assign unique ID to each semantic class as frame-wise labels. Based on SAM and SBICAC, we can train the proposed AdvImmu in an unsupervised manner.
The main contributions are summarized as follows:
  • This work proposes AdvImmu that is independent of reference images and resilient to various adverse weather conditions. It considers both local temporal correlations to enhance model performance and global randomness to prevent overfitting to specific temporal patterns.
  • AdvImmu also includes an image-level regularizer and an inter-class contrastive regularizer to enhance the model’s across-weather generalization, and additionally unfolds both regularizers into layers by deep unfolding to avoid instability and suboptimal performance.
  • The training of AdvImmu needs to overcome SAM’s category-agnostic issue in the output. To this end, SBICAC algorithm is proposed to generate category IDs serving as training labels based on SAM’s output.
  • Extensive experiments demonstrate that the proposed AdvImmu can generalize well under various adverse weather conditions. It outperforms existing state-of-the-art (SOTA) method by 88.56% in mean Intersection over Union (mIoU).
II Related Work
II-A AD Perception under Adverse Weather Conditions
AD perception under various adverse conditions has recently attracted increasing attention due to the strict safety demand in various outdoor navigation [9, 10] and perception tasks [11, 12, 13, 14]. Domain adaption [15] is the major scheme in enhancing AD perception capabilities under various adverse conditions. Some studies in domain adaption [6, 16] aim to learn a generalizable model from single or multiple related but distinct source domains where target data is inaccessible during model learning. Some adversarial training-based domain adaption works attempt to narrow the domain gap via style transfer or learning indistinguishable representations [17]. Furthermore, equipped with abundant data from various domains, curriculum learning-based domain adaption realize a progressive adaptation towards a distant target domain [18]. In addition, extending domain adaptation to multiple target domains usually employs domain transfer [19] to enhance generalization across domains. In this work, unlike domain adaption, the proposed AdvImmu can sense instantaneous, stable backgound, and dynamic information by considering temporal correlations, which is robust to arbitrary adverse conditions.
II-B Large Models (LMs) in AD
Large Models (LMs) have gained significant interest in the context of AD due to their proficiency in analyzing images and contexts [20]. These advancements have greatly improved zero-shot and few-shot image classification [21], segmentation [22, 23], object detection [24], etc. For example, CLIP [25] have shown that training to match images with captions can effectively create image representations from scratch; LLaMa [26] combines a vision encoder with an Large Language Model to enhance the understanding of both visual and linguistic concepts. Furthermore, researchers have explored the vectorized visual embeddings to equip LMs with environmental perception capabilities, particularly in AD scenarios. For example, DriveGPT4 [27] interprets video inputs to generate driving-related textual responses; Talk2BEV [28] leverages pre-trained image-language models to combine Bird’s Eye View maps with linguistic context, enabling visuo-linguistic reasoning in autonomous vehicles. In addition, generative models in AD are also gaining popularity. For instance, GAIA-1 [29] generates realistic driving scenarios by integrating video, text, and action inputs and showcases generative models in adapting to the changing dynamics of the real world. In this work, we propose to use SAM [8] to generate frame-wise annotation. However, SAM’s output lacks of category information. To mitigate this limitation, we additionally propose SBICAC clustering algorithm to supplement SAM to generate category IDs.
III Methodology
To depict the proposed AdvImmu clearly, we will introduce LSM and GSM together and denote the combination of them as LSaGS hereafter. Subsequently, we will detail URs and how the proposed SBICAC algorithm along with SAM to generate pseudo-labels.
III-A LSaGS
LSM extracts immediate, background, and dynamic features through InsU, IntU, and DU, respectively. Concretely, InsU is designed to process the current image as vehicle drives. It extracts instantaneous information from this image, which is crucial for the vehicle to accurately understand immediate surroundings. IntU focuses on processing stable background information from the environment, achieved by analyzing past consecutive images prior to the current frame. This enables IntU to understand what remains constant over time, which is important for consistent understanding of the scene. DU analyzes dynamic changes in the environment by comparing the current frame against previous ones, which helps to highlight the dynamic objects, such as pedestrians and vehicles. Notably, LSM depth is used to indicate the count of past consecutive frames that are considered.
Early Fusion (EF) and Late Fusion (LF) are two policies for fusing features from InsU, IntU, and DU. In EF strategy, features from InsU, IntU, and DU are first merged. Then the merged features are processed by a shared Convolution Neural Network (CNN) model. This early integration allows the model to utilize the correlations between these different types of features, enhancing the model for understanding street scenes. On the other hand, in LF strategy, InsU, IntU, and DU process the their own features by three separate CNN models. This allows each type of features to be finely analyzed on its own. After processing, these different types of features are combined. EF is summarized in the right part of Fig. 1 and LF is outlined in Fig. 2.
GSM is introduced to shuffle and group different LSM segments into mini-batches. In this way, each mini-batch reflects the diversity of the dataset instead of following a strict sequence. This enables the model learn more general and robust patterns, enhances its performance and generalization. Based on the shuffled mini-batches by GSM, the involved CNN(s) and the downstream perception head are then optimized by back propagation.
In summary, LSaGS is finally outlined in Algo. 1.
Figure 2: Illustration of LF.
Input: Input sequence $\{F\}$, LSM depth $d$, Fusion method $FusionType$, Epochs $epochs$
Output: $\text{Head},\text{CNN}_{\text{shd}}\ (or\ \text{CNN}_{\text{InsU/IntU/DU}})$
1 for $i\leftarrow 1$ to $epochs$ do
       // LSM
2       for frame $F_{t}$ in $\{F\}$ do
             $InsU_{t}\leftarrow\phi_{\text{InsU}}(F_{t})$ ;
              // Process immediate features from $F_{t}$
3
            $IntU_{t}\leftarrow\phi_{\text{IntU}}(\{F_{t-i}\}_{i=1}^{d})$ ;
              // Analyze background from past $d$ frames
4
            $DU_{t}\leftarrow\phi_{\text{DU}}(\{F_{t}-F_{t-i}\}_{i=1}^{d})$ ;
              // Analyze changes of $F_{t}$ to past $d$ frames
5
            // Fusion of InsU, IntU, and DU
6             if $FusionType==\text{EF}$ then
7                   $MF_{t}\leftarrow\text{Merge}(InsU_{t},IntU_{t},DU_{t})$
8                   $PF_{t}\leftarrow\text{CNN}_{\text{shd}}(MF_{t})$
9
10             end if
11            else if $FusionType==\text{LF}$ then
12                   $PInsU_{t}\leftarrow\text{CNN}_{\text{InsU}}(InsU_{t})$
13                   $PIntU_{t}\leftarrow\text{CNN}_{\text{IntU}}(IntU_{t})$
14                   $PDU_{t}\leftarrow\text{CNN}_{\text{DU}}(DU_{t})$
15                   $PF_{t}\leftarrow\text{Merge}(PInsU_{t},PIntU_{t},PDU_{t})$
16
17             end if
18
19       end for
      // GSM
20       $MiniBatches\leftarrow\text{Shuffle\&Group }\{PF_{t}\}$
21       for mini-batch $B$ in $MiniBatches$ do
22             $\text{Head},\text{CNN}_{\text{shd}}\ (or\ \text{CNN}_{\text{InsU/IntU/DU}})\leftarrow B$
23
24       end for
25
26 end for
Algorithm 1 LSaGS
III-B Unfolded Regularization
On top of the conventional pixel-level cross entropy loss (denoted as $L_{CE}$) in semantic understanding, we propose an image-level regularizer and an inter-class contrastive regularizer to enhance the model’s across-weather generalization. URs are the deep unfolding implementation of such regularizers, which aims at avoiding the heuristics and exhaustive searches on weights of them.
III-B1 The proposed Regularizers
We denote the distrituion of predicted masks $X$ and the ground truth $Y$ as $P_{X}$ and $P_{Y}$, respectively. We propose using Bhattacharyya distance (BD) as the image-level regularizer to quantify the distribution similarity between $X$ and $Y$. The definition of the BD regularizer over pixel $x$ is given by:
$L_{BD}=D_{B}(P_{X},P_{Y})=-\ln(\sum\nolimits_{x}\sqrt{P_{X}(x)P_{Y}(x)}),$ (1)
On the other hand, we propose using InfoNCE as an inter-class contrastive regularizer that can pull pixles of same class closer while push pixles of different classes apart. The InfoNCE regularizer is expressed as:
$L_{con}\!=\!-\!\log\!\frac{\sum_{m=1}^{|\mathcal{U}|}\exp(\frac{\sigma(x,u_{m})}{\tau})}{\sum_{m=1}^{|\mathcal{U}|}\!\exp(\frac{\sigma(x,u_{m})}{\tau})\!+\!\sum_{n=1}^{|\mathcal{V}|}\exp(\frac{\sigma(x,v_{n})}{\tau})},$ (2)
where $x$ is an anchor pixel, $u_{m}$ is any pixel from the same-class set $\mathcal{U}$, $v_{n}$ is any pixel from the different-class set $\mathcal{V}$, $\sigma(\cdot,\cdot)$ denotes the inner product, and $\tau$ controls the separation sharpness.
Figure 3: Illustration of URs. (4), (5), (6) represent equation number.
Figure 4: Overview of SAM-aided auto-annotation system. SAM and SBICAC in the dashed box just operate in training phase.
III-B2 Unfolded Regularizers (URs)
The total loss $L$ for training is:
$L=L_{CE}+\alpha L_{BD}+\gamma L_{con},$ (3)
where $\alpha$ and $\gamma$ are the weights assigned to $L_{BD}$ and $L_{con}$, respectively. In practice, they are tuned based on heuristics or exhaustive searches, which results in instability and suboptimal performance. Treating $\alpha$ and $\beta$ as trainable parameters can avoid these issues.
To this end, inspired by [7], we propose to treat each optimization iteration as one layer, and treat $\alpha$ and $\beta$ as learnable parameters within each layer. Based on this, above both regularizers can be unfolded into $K$ layers (illustrated in Fig. 3). Specifically, for layer $k,\text{where}\ k\in\{1,2,\cdots,K\}$, the output is updated as follow:
$x^{(k)}=\hskip 4.26773pt$ $x^{(k-1)}-\eta^{(k)}(\alpha^{(k)}\nabla_{x}L_{BD}(x^{(k-1)})\hskip 4.26773pt+\hskip 4.26773pt$
$\gamma^{(k)}\nabla_{x}L_{con}(x^{(k-1)})),$ (4)
where $\alpha^{(k)}$ and $\gamma^{(k)}$ are the learnable weights, $\eta^{(k)}$ is the step size of update, $\nabla_{x}L_{BD}(x^{(k-1)})$ and $\nabla_{x}L_{con}(x^{(k-1)})$ are the gradients of $L_{BD}$ and $L_{con}$ relative to pixel $x^{(k-1)}$, respectively, and they can be formulated as follows (denoting $x^{(k-1)}$ as $x$ for short):
$\nabla_{x}L_{BD}(x)$ $=-\frac{1}{2}\frac{1}{\sum_{x}\sqrt{P_{X}(x)P_{Y}(x)}}\frac{\sqrt{P_{Y}(x)}}{\sqrt{P_{X}(x)}},$ (5)
$\nabla_{x}L_{\text{con}}(x)$ $=\nabla_{x}B/B-\nabla_{x}A/A,\vspace{-0.9cm}$ (6)
where
$A$ $=\sum\nolimits_{m=1}^{|\mathcal{U}|}\exp(\sigma(x,u_{m})/\tau),$ (7)
$B$ $=A+\sum\nolimits_{n=1}^{|\mathcal{V}|}\exp(\sigma(x,v_{n})/\tau),$ (8)
$\nabla_{x}A$ $=\frac{1}{\tau}\sum\nolimits_{m=1}^{|\mathcal{U}|}\exp(\sigma(x,u_{m})/\tau)\cdot u_{m},$ (9)
$\!\nabla_{x}B\!$ $=\!\nabla_{x}A\!+\!\frac{1}{\tau}\sum\nolimits_{n=1}^{|\mathcal{V}|}\!\exp(\sigma(x,v_{n})/\tau)\cdot v_{n}.$ (10)
After multiple gradient descent updates by $K$ sequentially-connected layers, the output of the $K$-th layer is optimized by both regurlarizers, and then is added to the output of LSaGS. Then we use the summation results serving as predicted mask and the ground truth to calculate the cross entropy loss to optimize LSaGS and $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}$ by back propagation. In conclusion, URs is outlined in Algo. 2.
Input: $X$ (Input image), $Y$ (One-hot ground truth), $K$ (Layer number)
Output: Model $LSaGS$, Learnable variables $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}$
1 Initialize model $LSaGS\leftarrow\mathcal{W}_{0}$, and $\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)}\leftarrow\alpha_{0},\gamma_{0},\eta_{0},\text{where}\ k\in\{1,2,\cdots K\}$
2 for $epoch\ i\leftarrow 1$ to $max\_epochs$ do
3       $x^{(0)}\leftarrow LSaGS(X)$
       // Unfolding layers
4       for $layer\ k\leftarrow 1$ to $K$ do
             // Compute gradients
[…]
             // Update rule
[…]
8
9       end for
10      $x\leftarrow x^{(0)}+x^{(K)}$
11       $L=CrossEntropy(x,Y)$
12       $LSaGS,\alpha^{(k)}$, $\gamma^{(k)}$, $\eta^{(k)},\text{where}\ k\in\{1,2,\cdots K\}\leftarrow L.Backward()$
13 end for
Algorithm 2 Unfolded Regularizers (URs)
III-C SAM-Aided Consecutive Frame Annotation
SAM, a vision foundation model known for its strong zero-shot inference abilities, is used to generate masks for images in our research. It should be noted that SAM operates online during the training phase and offline during the inference phase, as illustrated in Fig. 4. However, SAM does not support generating category information (as demonstrated in LABEL:Fig.SAM_reason_to_cluster). To overcome this limitation, SBICAC algorithm is introduced to assign each semantic class a unique ID, which are then served as labels for training the model. SBICAC follows an iterative refinement manner and each iteration is described as follows:
  1. 1.
    Similarity Computation: For each pixel in $X$, the similarity to each cluster $centroid$ is computed using the inner product. The results store in a similarity matrix where each element represents the similarity between a pixel and a cluster $centroid$.
  2. 2.
    Cluster Assignment: Each pixel is assigned to the cluster with which it has the highest similarity. This is achieved by finding the index of maximum value in the similarity matrix for each pixel.
  3. 3.
    Cluster Centroid Calculation: The centroids are then updated based on the new cluster assignments. Specifically, for cluster $j$, the $new\_centroid$ is computed by averaging all pixel representation of that cluster:
    $new\_centroids[j]=\frac{1}{|X_{j}|}\sum X_{j},$
    where $X_{j}$ is the set of pixels assigned to cluster $j$.
  4. 4.
    Convergence Check and Centroid Update: SBICAC checks for convergence by comparing $new\_centroids$ to $centroids$. If $new\_centroids$ have not changed significantly (measured by a function like $np.allclose$), SBICAC terminates earlier; otherwise, $centroids$ are updated by $new\_centroids$ and next iteration begins.
In summary, SBICAC leverages inner product similarity to iteratively assign pixels to clusters and update clusters’ centroids until the cluster assignments stabilize. LABEL:Fig.cluster_comp exemplifies the performance of SBICAC against other clustering algorithms. This method ensures that the clusters are formed based on maximizing the similarity between pixels and their respective centroids. Upon completion of the iterative process, SBICAC outputs the cluster assignments $labels$ for all pixels. SBICAC is outlined in Algo. 3.
Input: $X$ (Latent feature maps), $n\_clusters$ (Number of clusters), $init$ (Initial random centroids), $max\_iter$ (Maximum iterations)
Output: Cluster assignments $labels$
1 Initialize centroids as $centroids\leftarrow init$
2 for $i\leftarrow 1$ to $max\_iter$ do
3       Compute similarity between $X$ and $centroids$: $similarity\leftarrow X\cdot centroids^{T}$
4       Assign clusters based on maximum similarity: $labels\leftarrow\arg\max(similarity,axis=1)$
5       Cluster Centroid Calculation:
6       for $j\leftarrow 1$ to $n\_clusters$ do
7             $cluster\ j$: $new\_centroids[j]\leftarrow\frac{1}{|X_{j}|}\sum X_{j}$
8       end for
9      Check for convergence:
10       if np.allclose$(centroids,new\_centroids)$ then
11             break
12
13       end if
14      Update centroids: $centroids\leftarrow new\_centroids$
15
16 end for
Algorithm 3 Similarity-based Iterative Cluster Assignments and Centroids Refinement (SBICAC)
IV Experiments
IV-A Datasets, Metrics and Implementation
IV-A1 Datasets
In our research, we utilize two datasets: the Apolloscapes dataset [30] and the CARLA_ADV dataset derived from the Carla simulator (version 0.9.13) [31]. The Apolloscape dataset is a comprehensive AD dataset featuring a variety of driving scenarios captured under various weather conditions. Each annotated image has pixel-level labels across 23 classes such as vehicles and pedestrians. The CARLA_ADV dataset is specifically designed to encompass a range of adverse weather conditions including fog, clouds, rain, darkness, and combinations thereof. It does also contain 23 semantic classes in pixel-level, including vehicle, building, tree, etc.
IV-A2 Metrics
We evaluate the proposed AdvImmu on street scene understanding task by employing four metrics: Mean Intersection over Union (mIoU), which measures the overlap between predicted mask and ground truth; Mean Precision (mPre), which assesses the accuracy of positive predictions; Mean Recall (mRec), evaluating how well the model identifies all relevant instances; and Mean F1 (mF1), which provides a balance between precision and recall.
TABLE I: The performance comparison of AdvImmu against other baselines
Methods Apolloscape CARLA_ADV
mIoU mPre mRec mF1 mIoU mPre mRec mF1
DeepLabv3+ [32] 26.58$\pm$0.49 30.23$\pm$0.89 31.14$\pm$0.57 32.32$\pm$0.59 36.90$\pm$1.00 46.96$\pm$0.97 42.44$\pm$0.79 43.78$\pm$0.93
BiSeNetV2 [33] 22.92$\pm$0.86 27.85$\pm$1.20 27.10$\pm$1.11 27.12$\pm$1.57 28.80$\pm$1.93 34.47$\pm$2.23 33.46$\pm$2.19 33.59$\pm$2.30
SegNet [34] 21.01$\pm$0.51 24.80$\pm$1.11 24.96$\pm$0.54 24.60$\pm$0.70 31.67$\pm$2.11 38.45$\pm$3.09 36.55$\pm$2.44 37.15$\pm$2.70
AdvImmu (Ours) 59.35$\pm$0.57 81.38$\pm$2.22 60.76$\pm$0.49 65.28$\pm$0.55 69.58$\pm$1.79 85.71$\pm$2.01 70.78$\pm$1.56 75.01$\pm$1.56
IV-A3 Implementation Details
Our deep learning model was implemented using the Pytorch framework, with experiments conducted on two NVIDIA GeForce 4090 GPUs. For the architecture, we implement the proposed AdvImmu based on ResNet [35] and ASSP perception head [32]. We select the Adam optimizer for training, configuring it with Betas values of 0.9 and 0.999, and set the weight decay at 1e-4. The training was executed with a batch size of 8 and a learning rate of 3e-4.
IV-B AdvImmu’s Overall Performance Evaluation
Tab. I compares the performance of the AdvImmu against DeepLabv3+ [32], BiSeNetV2 [33] and SegNet [34]. Across all metrics, AdvImmu generally outperforms DeepLabv3+, BiSeNetV2 and SegNet on both datasets. For example, on Apolloscapes dataset, AdvImmue outperforms DeepLabv3+, BiSeNetV2 and SegNet by (59.35 - 26.58) / 26.58 = 123.29%, (59.35 - 22.92) / 22.92 = 158.94% and (59.35 - 21.01) / 21.01 = 182.48% in mIoU, respectively; on CARLA_ADV dataset, AdvImmue outperforms DeepLabv3+, BiSeNetV2 and SegNet by (69.58 - 36.90) / 36.90 = 88.56%, (69.58 - 28.80) / 28.80 = 141.60% and (69.58 - 31.67) / 31.67 = 119.70% in mIoU, respectively. This suggests that AdvImmu should be more effective compared to the SOTA DeepLabv3+ model.
LABEL:Fig.perf_comp_inf compares the inference performance of AdvImmu against DeepLabv3+ (performing best in three baselines) in real driving test. Notably, AdvImmu consistently demonstrates better performance relative to DeepLabv3+, which can be supported by following aspects: (I) AdvImmu consistently achieves higher scores especially in mF1 and mIoU across all sequential frames, indicating better accuracy. (II) AdvImmu shows small performance variance across varied test conditions, such as fog, cloudy, rainy, dark, and combinations thereof, suggesting robustness against adverse weather conditions.
TABLE II: The effect of LSM depth on AdvImmu performance
LSM Depth mIoU mPre mRec mF1
Depth-1 41.52$\pm$0.77 57.83$\pm$0.92 46.90$\pm$0.67 49.96$\pm$0.78
Depth-2 41.29$\pm$0.80 57.43$\pm$0.60 46.68$\pm$0.84 49.76$\pm$0.80
Depth-3 41.58$\pm$0.78 57.52$\pm$0.85 47.08$\pm$0.72 50.09$\pm$0.75
Depth-4 41.81$\pm$0.88 56.78$\pm$0.70 46.23$\pm$0.81 49.13$\pm$0.84
IV-C In-depth Insights of AdvImmu
IV-C1 LSM Depth on Performance
Tab. II explores the effect of the LSM depth on performance. Across different LSM depths (Depth-1 through Depth-4), the performance metrics show certain variation, indicating that changes in depth have a little impact on AdvImmu’s performance. LABEL:Fig.LSM_depth_comp corroborates this finding, showing that all depths reach different performance levels, with slight fluctuations. However, the subtle variations in performance across different depths indicate that both too shallow (Depth-1) and too deep (Depth-4) configurations do not consistently yield the best results. Depths-3 generally show more stable and slightly improved performances, suggesting that moderate depths could be optimal. This implies that an intermediate LSM depth could potentially offer a better compromise between computation complexity and performance.
IV-C2 EF vs LF
LABEL:Fig.EF_vs_LF presents the comparison between EF and LF. LABEL:Fig.EF_arch and LABEL:Fig.LF_arch illustrate the architecture of EF and LF of AdvImmu. LABEL:Fig.EF_vs_LF_perf compares the performance of EF and LF across four metrics, where LF performs slightly better in mIoU and mPre, while EF has a slight advantage in mRec and mF1. Overall, the differences between the two fusion methods are minimal, suggesting that both approaches offer comparable performance with slight variations depending on the specific metric being evaluated.
IV-C3 Enabling GSM vs Disabling GSM
Tab. III demonstrates that enabling GSM significantly improves the model performance across all measured metrics. LABEL:Fig.GSM_comp shows that with GSM enabled, the performance metrics are consistently higher and more stable across 200 epochs, compared to the lower and more fluctuating metrics when GSM is disabled. Overall, enabling GSM leads to better and more reliable performance.
TABLE III: The effect of GSM on AdvImmu performance
Loss mIoU mPre mRec mF1
Disabling GSM 27.12$\pm$1.84 46.81$\pm$1.76 32.96$\pm$1.69 35.36$\pm$1.94
Enabling GSM 41.52$\pm$0.77 57.83$\pm$0.92 46.90$\pm$0.67 49.96$\pm$0.78
TABLE IV: The regularizer comparison
Loss mIoU mPre mRec mF1
CE 26.77$\pm$0.77 38.35$\pm$0.78 31.09$\pm$0.64 32.54$\pm$0.87
VRs 31.49$\pm$0.34 50.23$\pm$0.89 35.56$\pm$0.40 37.84$\pm$0.43
URs ($L=2$) 52.13$\pm$2.94 68.06$\pm$1.21 53.96$\pm$2.86 58.02$\pm$2.57
URs ($L=5$) 59.35$\pm$0.57 81.38$\pm$2.22 60.76$\pm$0.49 65.28$\pm$0.55
IV-C4 Vanilla regurlarizers (VRs) vs URs
Tab. IV and LABEL:Fig.regularizer_comp compare the performance of the following cases: (I) only cross entropy loss (denoted as CE); (II) cross entropy plus vanilla regularizers (denoted as VRs); (III) cross entropy plus 2-layer unfolded regularizers (denoted as URs ($L=2$)); (IV) cross entropy plus 5-layer unfolded regularizers (denoted as URs ($L=5$)). By comparing these cases, we can observe the following patterns: (I) Tab. IV shows a consistent improvement in performance when additional regularization techniques are added to the basic CE loss, leading to better scores in mIoU, mPre, mRec, and mF1. (II) Tab. IV also demonstrates that URs outperform corresponding VRs in almost all metrics. For example, URs ($L=2$) and URs ($L=5$) achieve mIoUs of 52.13 and 59.35, respectively, compared to 31.49 achieved by VRs. (III) Tab. IV additionally shows that the more layers the regularizers are unfolded, the better the performance obtained. For instance, URs ($L=5$) achieve an mIoU of 59.35, better than the 52.13 achieved by URs ($L=2$). These patterns can be visually confirmed in LABEL:Fig.regularizer_comp. In addition, LABEL:Fig.regularizer_comp illustrates that URs ($L=5$) converges faster than URs ($L=2$).
IV-D SBICAC Evaluation
LABEL:Fig.cluster_comp presents a qualitative comparison between SBICAC and other clustering algorithms. To quantitatively evaluate the proposed SBICAC against such baselines, we compare the mIoU between the ground truth and the respective masks generated by such baselines. However, to overcome the issue of the mismatch of cluster IDs, an approach is structured as follows:
  1. 1.
    Enumeration of ID Mappings: The approach initiates by enumerating all possible mappings of cluster IDs between ground truth and the generated masks. This ensures that every potential permutation of cluster assignments is considered.
  2. 2.
    Computation of IoU: For each hypothesized mapping, the IoU metric is calculated.
  3. 3.
    Aggregation of IoU: To derive a correspondence, the highest IoU values for each class are identified, and their average is computed, resulting in the mIoU. This mIoU measures the overall alignment between the clustering output and the ground truth.
LABEL:Fig.clutering_perf_mIoU shows a quantitative comparison of four clustering algorithms. The proposed SBICAC algorithm performs the best with an mIoU of 57.00%. It is followed by SpectralClustering at 53.10%, AgglomerativeClustering at 46.80%, and KMeans with the lowest mIoU of 30.40%. This demonstrates that SBICAC outperforms the other three algorithms. On the other hand, we utilize both the ground truth and SBICAC-generated pseudo-labels to train the proposed AdvImmu. As shown in LABEL:Fig.clutering_perf_training, the validation results in mIoU reveal that training with SBICAC-generated labels can achieve performance comparable to, or even surpassing, that of the ground truth. This superior performance is attributed to the fact that the generated labels are coarser than the ground truth (i.e., easier to classify).
V Conclusion
AD perception in various adverse weather conditions is challenging. This work proposed AdvImmu to enhance the model generalization across multiple adverse weather conditions by considering local temporal correlation. It processed frame sequences using InsU, IntU and DU to capture key information, consistent background and dynamic changes, and then shuffles these processed features to avoid overfitting and improve generalization. The training of AdvImmu requires frame annotations, not usually available in AD. To overcome this, we integrated the SAM and SBICAC clustering algorithm to support training. Our tests show that AdvImmu outperforms existing methods significantly.